# Mybibtex.bib

### v4

@@ Abstract
Existing Incremental Object Detection (IOD) methods partially alleviate catastrophic forgetting when incrementally detecting new objects in real-world scenarios. However, many of these methods rely on the assumption that unlabeled old-class objects may co-occur with labeled new-class objects in the incremental data. When unlabeled old-class objects are absent, the performance of existing methods tends to degrade. The absence can be mitigated by generating old-class samples, but it incurs high costs. This paper argues that previous generation-based IOD suffers from redundancy, both in the use of generative models, which require additional training and storage, and in the overproduction of generated samples, many of which do not contribute significantly to performance improvements. To eliminate the redundancy, we propose Inversed Objects Replay (IOR). Specifically, we generate old-class samples by inversing the original detectors, thus eliminating the necessity of training and storing additional generative models. We propose augmented replay to reuse the objects in generated samples, reducing redundant generations. Moreover, we propose high-value knowledge distillation focusing on the positions of old-class objects overwhelmed by the background, which transfers the knowledge to the incremental detector. Extensive experiments conducted on MS COCO 2017 demonstrate that our method can efficiently improve detection performance in IOD scenarios with the absence of old-class objects.
@@ II-B Inverse Detector Generation
where ${\mathcal{L}_{detect}}$ is the loss function between the prediction of the original detector and the sampled label $Y_{inv}$. ${\mathcal{L}_{detect}}$ is the same as the detection loss of the selected detector and is responsible for determining the category and position of the generated object in ${I_{inv}}$. The sampled label ${Y_{inv}}$ consists of five parameters, an object class $cls$ and four object bounding coordinates $x$, $y$, $w$, and $h$.
To sample a more realistic label ${Y_{inv}}$, we additionally preserve the histograms of the real object’s aspect ratios of each class. When inversing the detector, each image ${I_{inv}}$ is assigned a sampled label ${Y_{inv}}$. In label ${Y_{inv}}$, $cls$ is the desired class; the position $x$, $y$, $w$ is determined by a uniform distribution; the position $h=w\times ratio$, where $ratio$ is sampling from the preserved histogram. In this way, we can ensure that the generated object has a consistent aspect ratio with the real object, thus generating the object more accurately.
@@ II-C Augmented Replay
where mixing coefficient $\lambda$ is in the range $[0,1]$ and is sampled from the Beta distribution; $b(\hat{x},\hat{y})$ is a cropped bounding box with a randomly assigned location; $g\cup b$ is the intersection over union (IOU) between each $g$ and $b$; $thr$ is a threshold value. If the maximum IOU between each $g$ and $b$ is less than or equal to $thr$, then the original pixel value $\hat{I}_{inc}(x,y)$ is a mixture of the pixel value $I_{inc}(x,y)$ and the corresponding pixel value in the cropped bounding box $b$. If the maximum IOU between a given $g$ and $b$ is greater than $thr$, then the location of the $b(\hat{x},\hat{y})$ requires reassignment. To reuse the generated objects better, multiple instances of $b$ can be replayed within an image $I_{inc}$, forming the set of replayed bounding boxes $B_{r}$.

### v5

@@ Abstract
Existing Incremental Object Detection (IOD) methods partially alleviate catastrophic forgetting when incrementally detecting new objects in real-world scenarios. However, many of these methods rely on the assumption that unlabeled old-class objects may co-occur with labeled new-class objects in the incremental data. When unlabeled old-class objects are absent, the performance of existing methods tends to degrade. The absence can be mitigated by generating old-class samples, but it incurs high costs. This paper argues that previous generation-based IOD suffers from redundancy, both in the use of generative models, which require additional training and storage, and in the overproduction of generated samples, many of which do not contribute significantly to performance improvements. To eliminate the redundancy, we propose Inversed Objects Replay (IOR). Specifically, we generate old-class samples by inversing the original detectors, thus eliminating the necessity of training and storing additional generative models. We propose augmented replay to reuse the objects in generated samples, reducing redundant generations. Moreover, we propose high-value knowledge distillation focusing on the positions of old-class objects overwhelmed by the background, which transfers the knowledge to the incremental detector. Extensive experiments conducted on MS COCO 2017 demonstrate that our method can efficiently improve detection performance in IOD scenarios with the absence of old-class objects. The code is available at https://github.com/JiaJia075/IOR.
@@ II-B Inverse Detector Generation
where $\mathcal{L}_{detect}$} is the loss function between the prediction of the original detector and the sampled label $Y_{inv}$. $\mathcal{L}_{detect}$} is the same as the detection loss of the selected detector and is responsible for determining the category and position of the generated object in $I_{inv}$}. The sampled label $Y_{inv}$} consists of five parameters, an object class clscls and four object bounding coordinates $x$, $y$, $w$, and $h$.
To sample a more realistic label $Y_{inv}$}, we additionally preserve the histograms of the real object’s aspect ratios of each class. When inversing the detector, each image $I_{inv}$} is assigned a sampled label $Y_{inv}$}. In label $Y_{inv}$}, clscls is the desired class; the position $x$, $y$, $w$ is determined by a uniform distribution; the position $h=w\times ratio$, where ratioratio is sampling from the preserved histogram. In this way, we can ensure that the generated object has a consistent aspect ratio with the real object, thus generating the object more accurately.
@@ II-C Augmented Replay
[…]